OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work

OpenStudy (anonymous):

C/3 = 2f - 5h C/3 + 5h = 2f (C/3 + 5h)/2 = f C/6 + 2.5h = f

OpenStudy (anonymous):

C = 3(2f - 5h) C/3 = 2f - 5h (C/3) + 5h = 2f ((C/3) + 5h )/ 2 = f Therefore, f = (C/6) + 2.5h

OpenStudy (anonymous):

\[\frac{C}{3}=2f-5h\] \[\frac{C}{3}+5h=2f\] \[f=\frac{C}{6}+\frac{5h}{2}\]